# Emergence of Fermi’s Golden Rule in the Probing of a Quantum Many-Body System

### v1

@@ I Density homogeneity of the initial and outcoupled internal states
The realization of uniform quantum gases confined in optical boxes greatly simplifies the interpretation of density-dependent phenomena and eliminates inhomogeneous broadening in spectroscopy [47, 17]. Here we characterize the density homogeneity of the initial spin mixture and of the outcoupled state. Our system initially consists of a two-component gas of 6Li fermions in a spin-balanced mixture. The atoms are trapped in a cylindrical optical box trap (for more technical details, see e.g. [48, 33]). Examples of in situ absorption images of the occupied internal state $\ket{\downarrow}$ before the spectroscopic probe and the final outcoupled state |$\ket{\text{o}}$ after the spectroscopic probe are shown on the left of Fig. S1. The density profiles correspond respectively to the cuts along the dashed line on the OD images (the middle and rightmost profiles are respectively the cuts along the radial and axial directions). These profiles agree very well with fits that assume a homogeneous density inside the box (see solid lines).
@@
Figure S1: Spatial homogeneity of the initial spin-mixture and of the outcoupled state in the optical box. The upper (resp. lower) row corresponds to the state $\ket{\downarrow}$ before the spectroscopic probe (resp. |$\ket{\text{o}}$ after the spectroscopic probe). The color code is the same as Fig. 1 of the main text. The solid lines on the density profile cuts are fits assuming a uniform density distribution within a box trap. The vertical dotted lines mark the edge of the box trap and are located at the same position for each column. Here $N_{\downarrow}=1.71(5)\times 10^{5}$ ($\approx N_{\uparrow}$) and $N_{\text{o}}=2.2(1)\times 10^{4}$.
II Validity range of FGR away from the resonance
@@
The validity range of FGR with respect to $\Omega_{0}$ is $\omega$ dependent and is most stringent at $\omega=\omega_{\mathrm{p}}$. In the main text, we observed the on-resonance power broadening of the spectral response. Here we briefly look into the applicability of FGR away from $\omega_{\mathrm{p}}$. In Fig. S2, we show the measurement of the linear transfer rate $I_{\text{linfit}}$ for a detuned drive of frequency $\hbar\omega/E_{\mathrm{F}}=0.76(2)$, which is approximately $\Gamma_{0}/2$ away from $\omega_{\text{p}}$ (see blue points in the inset of Fig. S2); we use pulses of duration up to $2\pi/(5\Omega_{0})$. We show this data alongside the original data of the lower panel of Fig. 4 (orange points). For the detuned drive, the response remains unchanged up to $\hbar\Omega_{0}/E_{\mathrm{F}}\approx 0.07$; in the resonant case, deviations can already be seen for $\hbar\Omega_{0}/E_{\mathrm{F}}\gtrsim 0.01$.
@@
Figure S2: Validity of FGR away from the resonance. The linear rate is extracted with a linear fit of the transfer fraction with pulse durations up to $t=2\pi/(5\Omega_{0})$. Inset: same spectrum as in the upper inset of Fig. 3, where color markers are adjusted to match the main panel, i.e. the blue (resp. orange) point corresponds to the detuned (resp. on-resonance) response. Note that for this spectrum, $\hbar\Gamma_{0}/E_{\mathrm{F}}\approx 0.24$.
@@ III Extraction of spectral properties
Figure S3: Robustness of spectral property extraction. (a) Example of an rf spectrum, together with two models for extracting the spectral properties. The solid blue (resp. dashed beige) line is a smoothened interpolation using an SG filter (resp. fitting with Lorentzian lineshape). (b) Comparison between the spectral properties extracted using the two models; the color style is the same as in (a).
In Fig. 4 of the main text, we extracted properties from the measured spectra using smoothened interpolations of the data via a Savitzky–Golay (SG) filter [49]. To validate the robustness of these extracted properties, we also use a Lorentzian fitting function. An example of a direct comparison on a typical spectrum is shown in Fig. S3a. In Fig. S3b, we also show the comparison of the extracted spectral properties. The parameters largely match (even though the Lorentzian fit performs badly in the tails of the spectrum).
@@ IV Calibration of $\Omega_{0}$ and Magnetic field stability
To provide an upper bound to our magnetic field stability, we show an rf spectrum in Fig. S4b taken for a pulse duration of $12~{}\mathrm{ms}$. The solid line is a fit to a sinc-squared function, whose FWHM is $2\pi\times 70(4)~{}\mathrm{Hz}$; this agrees well with the expected Fourier limit $\approx 2\pi\times 73.8~{}\mathrm{Hz}$. This implies that our magnetic field stability is better than $\approx 35~{}\mathrm{mG}$.
In Fig. 3, we restricted the data to pulse durations below $12~{}\mathrm{ms}$. In Figs. S4a and c, we show the data for pulse durations longer than $12~{}\mathrm{ms}$ as open markers. Although the transfer fraction is low, $\alpha<0.1$, we observe a deviation from the predicted solid black line. We attribute this to the magnetic field drifts over longer times.
Figure S4: Calibration of $\Omega_{0}$ and magnetic field stability. (a) Rabi oscillations of fully polarized gases using on-resonant rf. The solid lines are fits to the data. The dotted lines are the extension of the fits and the open markers are beyond the pulse duration where the Fourier-limited spectrum can be achieved. (b) Rf spectrum of a fully polarized gas that is Fourier limited for a pulse duration of $t=12~{}\mathrm{ms}$. The solid line is a fit using sinc-squared function. (c) Deviation from linear-response expectation due to magnetic field noise. The solid-marker data, the solid black line, and the dotted blue line are the same as in Fig. 3. Open markers are the measurements beyond our limit of $12~{}\mathrm{ms}$.
@@ V Linear response theory: FGR and beyond
In this section, we formally prove the different regimes that appeared in the linear response framework as depicted in Fig. 6. We use the lowest three hyperfine states of 6Li, with the general Hamiltonian
$\hat{H}=\hat{H}_{\mathrm{S}}+\hat{V}_{\mathrm{pr}}$ (S1)
where $\hat{H}_{\mathrm{S}}$ is the system Hamiltonian that includes the kinetic energy and the interaction energy, and
$\hat{V}_{\mathrm{pr}}=\frac{\hbar\Omega_{0}}{2}s(t)e^{-i\omega_{\mathrm{pr}}t}\sum_{\textbf{k}}\hat{c}_{\text{o},\textbf{k}}^{\dagger}\hat{c}_{\downarrow,\textbf{k}}+\mathrm{h.c.}$ (S2)
is the spectroscopic probe that couples state $\ket{\downarrow}$ and |$\ket{\text{o}}$ with a driving frequency $\omega_{\mathrm{pr}}$, where $\Omega_{0}$ is the bare-transition Rabi frequency, $s(t)$ is the temporal profile of the rf field and is assumed to be a square pulse in the main text for simplicity, $\hat{c}_{\sigma,\textbf{k}}$ ($\hat{c}_{\sigma,\textbf{k}}^{\dagger}$) is the annihilation (creation) operator for particles with state $\sigma$ and momentum $\hbar\textbf{k}$, and $\mathrm{h.c.}$ is the hermitian conjugate. We then use the transition rate as the observable, which is defined as
@@
$=-i\frac{\Omega_{0}}{2N_{\downarrow}}s(t)e^{-i\omega_{\mathrm{pr}}t}\sum_{\textbf{k}}\hat{c}_{\text{o},\textbf{k}}^{\dagger}\hat{c}_{\downarrow,\textbf{k}}+\mathrm{h.c.},$
@@ V.1 Linear response theory of the transfer rate
Using linear response theory and assuming |$\ket{\text{o}}$ is non-interacting with both $\ket{\downarrow}$ and $\ket{\uparrow}$, the time evolution of the transition rate in the interaction picture under rotating wave approximation, which is denoted by the superscript $\hat{R}^{I}(\omega,t)=e^{i\hat{H}_{\mathrm{S}}t/\hbar}\hat{R}(\omega,t)e^{-i\hat{H}_{\mathrm{S}}t/\hbar}$, is [50]
$R(\omega,t)\equiv\langle\hat{R}^{I}(\omega,t)\rangle$ $=-\frac{i}{\hbar}\int_{-\infty}^{\infty}\mathrm{d}t^{\prime}\Theta(t-t^{\prime})\Big{\langle}[\hat{R}^{I}(t),\hat{V}_{\mathrm{pr}}^{I}(t^{\prime})]\Big{\rangle}_{0}$ (S4)
$=\frac{\Omega_{0}^{2}}{4N_{\downarrow}}\sum_{k}\int_{0}^{t}\mathrm{d}t^{\prime}~{}e^{i(\xi_{k\downarrow}/\hbar-\omega)(t-t^{\prime})}(-i)G_{\downarrow}^{<}(\textbf{k},t-t^{\prime})+\mathrm{c.c.}$
$=\frac{\Omega_{0}^{2}}{4N_{\downarrow}}\sum_{k}\int_{0}^{t}\mathrm{d}t^{\prime}\int_{-\infty}^{\infty}\frac{\mathrm{d}\omega^{\prime}}{2\pi}~{}e^{i(\xi_{k\downarrow}/\hbar-\omega-\omega^{\prime})(t-t^{\prime})}A(\textbf{k},\omega^{\prime})n_{\mathrm{F}}(\hbar\omega^{\prime})+\mathrm{c.c.},$
where $\braket{\cdot}_{0}$ is the average over the equilibrium ensemble without the probe, $\omega=\omega_{\mathrm{pr}}-\omega_{0}$ is the detuning, $G_{\downarrow}^{<}(\textbf{k},t-t^{\prime})\equiv i\braket{\hat{c}_{\downarrow,\textbf{k}}^{\dagger}(t^{\prime})\hat{c}_{\downarrow,\textbf{k}}(t)}_{0}$ is the lesser Green’s function, $A(\textbf{k},\omega)$ is the spectral function related to the Fourier transform of the lesser Green’s function through $A(\textbf{k},\omega)n_{\mathrm{F}}(\hbar\omega)=G_{\downarrow}^{<}(\textbf{k},\omega)\equiv\int_{-\infty}^{\infty}e^{-i\omega t}G_{\downarrow}^{<}(\textbf{k},t)\mathrm{d}t$, and $n_{\mathrm{F}}(\epsilon)=1/(e^{(\epsilon-\mu)/k_{\mathrm{B}}T}+1)$ is the Fermi-Dirac distribution. A square pulse is assumed for the equation as well.
@@ V.2 Long-time limit: FGR
In the appropriate time range, FGR provides the simple expression for the rf spectrum $R_{\mathrm{FGR}}(\omega)$
$R_{\mathrm{FGR}}(\omega)\equiv\frac{\Omega_{0}^{2}}{4N_{\downarrow}}\sum_{\textbf{k}}A(\textbf{k},\xi_{k\downarrow}/\hbar-\omega)n_{\mathrm{F}}(\xi_{k\downarrow}-\hbar\omega),$ (S5)
where $\xi_{k\downarrow}=\hbar^{2}k^{2}/(2m)-\mu_{\downarrow}$ is the single-particle kinetic energy with $\textbf{k}\rvert$ k | and chemical potential $\mu_{\downarrow}$. The normalization of $R_{\mathrm{FGR}}(\omega)$ is
@@
In the main text, we defined a normalized spectrum $I_{\mathrm{FGR}}$ as $I_{\mathrm{FGR}}(\hbar\omega/E_{\mathrm{F}})=R_{\mathrm{FGR}}(\omega)\cdot(E_{\mathrm{F}}/\hbar\Omega_{0}^{2})$. This spectrum is normalized such that
$\int_{-\infty}^{\infty}\mathrm{d}(\hbar\omega/E_{\mathrm{F}})~{}I_{\mathrm{FGR}}(\hbar\omega/E_{\mathrm{F}})=\frac{\pi}{2}.$ (S7)
@@ V.3 Relation between linear response and FGR
We can obtain Eq. (2) for $\alpha$ via an integration:
@@
V.4 Short time limit: quadratic dependence in time
@@
Note that the sub-leading order can be calculated using the contact $C$, which characterizes the high-frequency limit of the rf spectrum [45, 51, 52]
$R_{\mathrm{FGR}}(\hbar\omega/E_{\mathrm{F}}\gg 1)\approx\frac{\sqrt{\hbar}\Omega_{0}^{2}}{8\pi N_{\downarrow}}\frac{C}{\sqrt{m}\omega^{3/2}}.$ (S12)
@@
where $t_{\mathrm{F}}\equiv h/E_{\mathrm{F}}$, and $\tilde{C}\equiv C/(N_{\downarrow}k_{\mathrm{F}})$.

### v2

@@ I Density homogeneity of the initial and outcoupled internal states
The realization of uniform quantum gases confined in optical boxes greatly simplifies the interpretation of density-dependent phenomena and eliminates inhomogeneous broadening in spectroscopy [48, 17]. Here we characterize the density homogeneity of the initial spin mixture and of the outcoupled state. Our system initially consists of a two-component gas of 6Li fermions in a spin-balanced mixture. The atoms are trapped in a cylindrical optical box trap (for more technical details, see e.g. [49, 33]). Examples of in situ absorption images of the occupied internal state $\ket{\downarrow}$ before the coupling is turned on and the final outcoupled state |$\ket{\text{o}}$ after the coupling pulse are shown on the left of Fig. S1. The density profiles correspond respectively to the cuts along the dashed line on the OD images (the middle and rightmost profiles are respectively the cuts along the radial and axial directions). These profiles agree very well with fits that assume a homogeneous density inside the box (see solid lines).
@@
Figure S1: Spatial homogeneity of the initial spin mixture and of the outcoupled state in the optical box. The upper (resp. lower) row corresponds to the state $\ket{\downarrow}$ before the coupling is turned on (resp. |$\ket{\text{o}}$ after the coupling pulse). The color code is the same as Fig. 1 of the main text. The solid lines on the density profile cuts are fits assuming a uniform density distribution within a box trap. The vertical dotted lines mark the edges of the box trap and are located at the same position for each column. Here $N_{\downarrow}=1.71(5)\times 10^{5}$ ($\approx N_{\uparrow}$) and $N_{\text{o}}=2.2(1)\times 10^{4}$.
II Validity range of FGR away from the peak response frequency
@@
The validity range of FGR with respect to $\Omega_{0}$ is $\omega$ dependent and is most stringent on resonance, $\omega=\omega_{\mathrm{p}}$. In the main text, we observed the on-resonance power broadening of the spectral response. Here we briefly look into the applicability of FGR away from $\omega_{\mathrm{p}}$. In Fig. S2, we show the measurement of the linear transfer rate $I_{\text{linfit}}$ for a detuned drive of frequency $\hbar\omega/E_{\mathrm{F}}=0.76(2)$, which is approximately $\Gamma_{0}/2$ away from $\omega_{\text{p}}$ (see blue points in the inset of Fig. S2); we use pulses of duration up to $2\pi/(5\Omega_{0})$. We show this data alongside the original data of the lower panel of Fig. 4 (light orange points). For the detuned drive, the response remains unchanged up to $\hbar\Omega_{0}/E_{\mathrm{F}}\approx 0.07$; in the resonant case, deviations can already be seen for $\hbar\Omega_{0}/E_{\mathrm{F}}\gtrsim 0.01$.
@@
Figure S2: Validity of FGR away from the resonance. The linear rate is extracted with a linear fit of the transfer fraction with pulse durations up to $t=2\pi/(5\Omega_{0})$. Inset: same spectrum as in the upper inset of Fig. 3, where color markers are adjusted to match the main panel, i.e. the blue (resp. light orange) point corresponds to the detuned (resp. on-resonance) response. Note that for this spectrum, $\hbar\Gamma_{0}/E_{\mathrm{F}}\approx 0.24$.
@@ III Extraction of spectral properties
Figure S3: Robustness of the extraction of the spectral property. (a) Example of an rf spectrum, together with two models for extracting the spectral properties. The solid blue (resp. dashed beige) line is a smoothened interpolation using an SG filter (resp. fitting with Lorentzian lineshape). (b) Comparison between the spectral properties extracted using the two models; the color style is the same as in (a).
In Fig. 4 of the main text, we extracted properties from the measured spectra using smoothened interpolations of the data via a Savitzky–Golay (SG) filter [50]. To verify the robustness of these extracted properties, we also use a Lorentzian fitting function. An example of a direct comparison on a typical spectrum is shown in Fig. S3a. In Fig. S3b, we also show the comparison of the extracted spectral properties. The parameters largely match (even though the Lorentzian fit performs badly in the tails of the spectrum).
@@ IV Calibration of $\Omega_{0}$ and Magnetic field stability
To provide a bound to our magnetic field stability, we measure long-pulse spectra; in Fig. S4b, we show a spectrum taken with a pulse duration of $12~{}\mathrm{ms}$. The solid line is a fit to a sinc-squared function, whose FWHM is $2\pi\times 70(4)~{}\mathrm{Hz}$; this agrees well with the expected Fourier limit $\approx 2\pi\times 73.8~{}\mathrm{Hz}$. This implies that our magnetic field stability is better than $\approx 35~{}\mathrm{mG}$.
As a result, we restricted in Fig. 3 the data to pulse durations below $12~{}\mathrm{ms}$. In Figs. S4a and c, we show the data for pulse durations longer than $12~{}\mathrm{ms}$ as open markers. Although the transfer fraction is low ($\alpha<0.1$), we observe a deviation from the predicted solid black line. We attribute this to magnetic field drifts over longer times.
Figure S4: Calibration of $\Omega_{0}$ and magnetic field stability. (a) Rabi oscillations in fully polarized gases using on-resonant rf. The solid lines are fits to the data. The dotted lines are the extension of the fits and the open markers are beyond the pulse duration where the Fourier-limited spectrum is achieved. (b) Rf spectrum of a fully polarized gas that is Fourier limited for a pulse duration of $t=12~{}\mathrm{ms}$. The solid line is a sinc-squared fit. (c) Deviation from linear-response expectation due to magnetic field noise. The solid-marker data, the solid black line, and the dotted blue line are the same as in Fig. 3. Open markers are the measurements beyond our limit of $12~{}\mathrm{ms}$.
@@ V Linear response theory: FGR and beyond
In this section, we theoretically outline the different regimes within the linear response region depicted in Fig. 6. The full Hamiltonian is
$\hat{H}(t)=\hat{H}_{\mathrm{S}}+\hat{V}_{\mathrm{pr}}(t)$ (S1)
where $\hat{H}_{\mathrm{S}}$ is the (time-independent) Hamiltonian of the spin-$1/2$ unitary Fermi gas, with spin states $\ket{\uparrow}$ and $\ket{\downarrow}$. The probing term that couples states $\ket{\downarrow}$ and |$\ket{\text{o}}$ is
$\hat{V}_{\mathrm{pr}}(t)=\hbar\Omega_{0}s(t)\cos\left(\omega_{\mathrm{pr}}t\right)\sum_{\textbf{k}}\hat{c}_{\text{o},\textbf{k}}^{\dagger}\hat{c}_{\downarrow,\textbf{k}}+\mathrm{h.c.},$ (S2)
where $\omega_{\mathrm{pr}}$ is the driving frequency, $\Omega_{0}$ is the bare-transition Rabi frequency, $s(t)$ is the temporal profile of the rf field (assumed to be a square pulse in the main text), and $\hat{c}_{\sigma,\textbf{k}}$ (resp. $\hat{c}_{\sigma,\textbf{k}}^{\dagger}$) is the annihilation (resp. creation) operator for particles with state $\sigma$ and momentum $\hbar\textbf{k}$. The transition rate into |$\ket{\text{o}}$ is our observable, defined as
@@
$=-i\frac{\Omega_{0}}{N_{\downarrow}}s(t)\cos\left(\omega_{\mathrm{pr}}t\right)\sum_{\textbf{k}}\hat{c}_{\text{o},\textbf{k}}^{\dagger}\hat{c}_{\downarrow,\textbf{k}}+\mathrm{h.c.},$
@@ V.1 Linear response theory of the transfer rate
We now calculate the time evolution of the transition rate as a function of drive detuning $\omega\equiv\omega_{\mathrm{pr}}-\omega_{0}$:
$R(\omega,t)\equiv\langle\hat{R}\rangle$ $\approx-\frac{i}{\hbar}\int_{-\infty}^{\infty}\mathrm{d}t^{\prime}\Theta(t-t^{\prime})\Big{\langle}[\hat{R}^{\mathrm{I}}(t),\hat{V}_{\mathrm{pr}}^{\mathrm{I}}(t^{\prime})]\Big{\rangle}_{0}$ (S4)
$\approx\frac{\Omega_{0}^{2}}{4N_{\downarrow}}\sum_{\textbf{k}}\int_{0}^{t}\mathrm{d}t^{\prime}~{}e^{i(\xi_{k\downarrow}/\hbar-\omega)(t-t^{\prime})}(-i)G_{\downarrow}^{<}(\textbf{k},t-t^{\prime})+\mathrm{c.c.}$
$=\frac{\Omega_{0}^{2}}{4N_{\downarrow}}\sum_{\textbf{k}}\int_{0}^{t}\mathrm{d}t^{\prime}\int_{-\infty}^{\infty}\frac{\mathrm{d}\omega^{\prime}}{2\pi}~{}e^{i(\xi_{k\downarrow}/\hbar-\omega-\omega^{\prime})(t-t^{\prime})}A(\textbf{k},\omega^{\prime})n_{\mathrm{F}}(\hbar\omega^{\prime})+\mathrm{c.c.},$
where $\hat{\mathcal{O}}^{\mathrm{I}}\equiv e^{i\hat{H}_{\mathrm{S}}t/\hbar}\hat{\mathcal{O}}e^{-i\hat{H}_{\mathrm{S}}t/\hbar}$ is the corresponding operator in the interaction picture. The approximations in the first and second lines are respectively linear response (i.e. we only keep the term to first order in $\Omega_{0}^{2}$) and rotating wave approximations. $\braket{\hat{\mathcal{O}}}_{0}\equiv\mathrm{Tr}[\hat{\rho}_{0}\hat{\mathcal{O}}]$ denotes the average over the initial equilibrium state $\hat{\rho}_{0}$ of a spin-balanced spin-1/2 unitary gas governed by the Hamiltonian $\hat{H}_{\mathrm{S}}$. $\braket{\hat{\mathcal{O}}}\equiv\mathrm{Tr}[\hat{\rho}(t)\hat{\mathcal{O}}]$ denotes the average over the state $\hat{\rho}(t)$ which at $t=0$ is the state $\hat{\rho}_{0}$ and which evolves according to the full time-dependent Hamiltonian Eq. (S1). We introduced the lesser Green’s function $G_{\downarrow}^{<}(\textbf{k},t-t^{\prime})\equiv ie^{i\mu_{\downarrow}(t-t^{\prime})}\braket{(\hat{c}_{\downarrow,\textbf{k}}^{\mathrm{I}}(t^{\prime}))^{\dagger}\hat{c}_{\downarrow,\textbf{k}}^{\mathrm{I}}(t)}_{0}$, and the single-particle kinetic energy $\xi_{k\downarrow}=\hbar^{2}k^{2}/(2m)-\mu_{\downarrow}$ with $\textbf{k}\rvert$ k | and chemical potential $\mu_{\downarrow}$. The spectral function $A(\textbf{k},\omega)$ is related to the Fourier transform of the lesser Green’s function through $A(\textbf{k},\omega)n_{\mathrm{F}}(\hbar\omega)=G_{\downarrow}^{<}(\textbf{k},\omega)\equiv\int_{-\infty}^{\infty}e^{i\omega t}G_{\downarrow}^{<}(\textbf{k},t)\mathrm{d}t$, where $n_{\mathrm{F}}(\epsilon)=1/(e^{{\beta}\epsilon}+1)$ is the Fermi-Dirac distribution, and $\beta^{-1}\equiv k_{\mathrm{B}}T$. In the second line, we assumed that the coupling is abruptly turned on at $t=0$.
@@ V.2 Long-time limit: FGR
In the long-time limit (while remaining within the validity regime of linear response theory), Eq. (S4) reduces to FGR, which provides a simple expression for the rf spectrum $R_{\mathrm{FGR}}(\omega)$:
$R_{\mathrm{FGR}}(\omega)\equiv\frac{\Omega_{0}^{2}}{4N_{\downarrow}}\sum_{\textbf{k}}A(\textbf{k},\xi_{k\downarrow}/\hbar-\omega)n_{\mathrm{F}}(\xi_{k\downarrow}-\hbar\omega).$ (S5)
The normalization of $R_{\mathrm{FGR}}(\omega)$ is
@@
In the main text, we defined a normalized spectrum $I_{\mathrm{FGR}}(\hbar\omega/E_{\mathrm{F}})\equiv R_{\mathrm{FGR}}(\omega)\cdot(E_{\mathrm{F}}/\hbar\Omega_{0}^{2})$. This spectrum is normalized such that
$\int_{-\infty}^{\infty}\mathrm{d}\left(\frac{\hbar\omega}{E_{\mathrm{F}}}\right)~{}I_{\mathrm{FGR}}\left(\frac{\hbar\omega}{E_{\mathrm{F}}}\right)=\frac{\pi}{2}.$ (S7)
@@ V.3 Relation between linear response and FGR
We obtain Eq. (2) from the main text via integration
@@
V.4 Short-time limit: quadratic dependence in time
@@
Note that the sub-leading order can be calculated using the contact $C$, which characterizes the high-frequency limit of the rf spectrum [46, 51, 52]:
$R_{\mathrm{FGR}}\left(\frac{\hbar\omega}{E_{\mathrm{F}}}\gg 1\right)\approx\frac{\sqrt{\hbar}\Omega_{0}^{2}}{8\pi N_{\downarrow}}\frac{C}{\sqrt{m}\omega^{3/2}}.$ (S12)
@@
where $t_{\mathrm{F}}\equiv\hbar/E_{\mathrm{F}}$, and $\tilde{C}\equiv C/(N_{\downarrow}k_{\mathrm{F}})$.